An electron has the following wave function:

ψ(x)=A(e^-x)(1-e^-x)


Determine
i) the normalization constant A, and
ii) the expectation value of x .
1
Expert's answer
2020-04-15T10:20:51-0400

The wave function of an electron

"\psi(x)=Ae^{-x}(1-e^{-x})."

(i) The normalization condition

"1=\int_{0}^{\infty}|\psi(x)|^2dx\\
1=A^2\int_{0}^{\infty}e^{-2x}(1-e^{-x})^2dx=\frac{A^2}{12}\\
A=2\sqrt{3}."

(ii) The expectation value of "x"

"\lang x\rang =12\int_{0}^{\infty}xe^{-2x}(1-e^{-x})^2dx=\frac{13}{12}=1.08"